OpenStudy (anonymous):

given f(x) = 3sin(2x) and g(x) = 3/2. evaluate algebraically the points where the 2 functions intercept domain of [0.2pi],

OpenStudy (anonymous):

does this really mean solve \[3\sin(2x)=\frac{3}{2}\]

OpenStudy (anonymous):

if so, then \[\sin(2x)=\frac{1}{2}\] which is easy enough to solve for \(2x\) then divide by 2 to get \(x\)

OpenStudy (anonymous):

that is what I did, is that correct?

OpenStudy (anonymous):

did you get \(2x=\frac{\pi}{6}\)?

OpenStudy (anonymous):

also \[2x=\frac{5\pi}{6}\]

OpenStudy (anonymous):

yes :) my next two I got 13pi/12 and 17pi/12

OpenStudy (anonymous):

looks reasonable

OpenStudy (anonymous):

cool, thank you :)

OpenStudy (anonymous):

yw